OpenStudy (anonymous):

what is the derivative of radical 2 over radical x.

OpenStudy (amistre64):

(2/x)^(1/2) (1/2)(2/x)^(-1/2) -2/x^2

OpenStudy (amistre64):

1 sqrt(x) -2 ---------- ? 2 sqrt(2) x^2

OpenStudy (anonymous):

\[\frac{\sqrt{2}}{\sqrt{x}}\]?

OpenStudy (amistre64):

- x^(1/2) -------- = -1/ x^(3/2) = -1/x sqrt(x) i think x^2

OpenStudy (anonymous):

umm..

OpenStudy (anonymous):

if so the answer is \[-\frac{\sqrt{2}}{2\sqrt{x^3}}\]

OpenStudy (anonymous):

okay yeah thats what I was looking for :D

OpenStudy (amistre64):

lol .... i was close ;)

OpenStudy (anonymous):

dont worry I gave both of you a medal :D

OpenStudy (anonymous):

I was doing taylor series and forgot how to find derivatives of those kinds of functions :\

OpenStudy (amistre64):

;)